# Quantum optical coherence theory based on Feynman’s path integral

### v1

@@ VI.1.2 First-order interference of two independent matter wave beams
Figure 34: Scheme for the first-order interference of two independent matter wave beams.
There are two different paths to trigger a particle detection event in the observation plane. The first one is the detected particle is emitted S1. The probability is 1/2 and the probability amplitude is $A_{1x}$. The second one is the detected particle is emitted S2. The probability is 1/2 and the probability amplitude is $A_{2x}$. If these two different paths are indistinguishable, the probability function in the observation plane is
$P^{(1)}(x,t)$ $=$ $\langle|A_{1x}+A_{2x}|^{2}\rangle$ (106)
$=$ $\langle|e^{i\varphi_{1j}}K_{1x}+e^{i\varphi_{2j}}K_{2x}|^{2}\rangle,$
where $\varphi_{1j}$ and $\varphi_{2j}$ are the initial phases of particles emitted by S1 and S2, respectively. With the same approximations above, spatial probability function can be obtained via Eq. (106),
@@
By analogy of the experimental results of the degree of second- and third-order coherence of cold atoms and BEC to thermal and laser light Yasuda and Shimizu (1996); Burt et al. (1997); Lopes et al. (2015), it is reasonable to assume the coherence properties of cold atoms are similar as the ones of thermal light and the coherence properties of BEC are similar as the one of laser light. The initial phases of particles in cold atoms are random. The initial phases of particles in one BEC are identical.
@@
where $N$ is the number of detected particles and $\varphi$ is a random phase due to the sum of $\varphi_{1j}-\varphi_{2j}$ for $N$ detected particles. The first-order interference pattern can not observed.
@@
$P^{(1)}(x)$ $=$ $1+\cos[\frac{2\pi d}{L\lambda_{D}}x+(\varphi_{1j}-\varphi_{2j})].$ (109)
$\varphi_{1j}-\varphi_{2j}$ is a random but fixed number for two independent BECs. The first-order interference pattern can be observed Andrews et al. (1997), just like the transient first-order interference pattern by superposing two independent laser light beams Magyar and Mandel (1963).
@@ VI.2.1 Second-order interference of one matter wave beam
The scheme for the second-order interference of one matter wave beam is similar as the one shown in Fig. 11 except the source S is a particle source, BS is a beam splitter for particles and these two detectors are particle detectors.
First discuss the second-order interference of one boson particle beam. Assuming S is a cold atom source, there are two different paths to trigger a two-particle coincidence count. The first one is particle a is detected by D1 and particle b is detected by D2. The probability is 1/2 and the corresponding two-particle amplitude is $A_{a1,b2}$. The second path is particle a is detected by D2 and particle b is detected by D1. The probability is 1/2 and the corresponding two-particle amplitude is $A_{a2,b1}$. If these two different paths are indistinguishable, the two-particle probability function is
@@
$=$ $\langle|A_{a1,b2}+A_{a2,b1}|^{2}\rangle$
$=$ $\langle|e^{i\varphi_{a}}K_{a1}e^{i\varphi_{b}}K_{b2}+e^{i\varphi_{a}}K_{a2}e^{i\varphi_{b}}K_{b1}|^{2}\rangle$
@@
where $D$ is the length of S and single speed particles are assumed. The temporal two-particle probability function is
@@
When these two detectors are at the same space-time coordinates, the probability to detect two particles is two times of the one when these two detectors are in different coherence volumes. There is two-particle bunching for cold boson particles in a HBT interferometer Yasuda and Shimizu (1996), just like there is two-photon bunching for thermal light in a HBT interferometer.
Assume S is a BEC in Fig. 11, there is only one path to trigger a two-particle coincidence count. Since the initial phases of particles in a BEC are identical, just like the one for single-mode continuous-wave laser light. No two-particle bunching can be observed, which is consistent with Burt et al.’s experiments Burt et al. (1997).
Then discuss the second-order interference of one fermion particle beam. There are also two different ways to trigger a two-particle coincidence count and the corresponding two-particle probability amplitude are $A_{a1,b2}$ and $A_{a2,b1}$. If these two different ways are indistinguishable, the two-particle probability function is
@@
$=$ $\langle|A_{a1,b2}-A_{a2,b1}|^{2}\rangle$
$=$ $\langle|e^{i\varphi_{a}}K_{a1}e^{i\varphi_{b}}K_{b2}-e^{i\varphi_{a}}K_{a2}e^{i\varphi_{b}}K_{b1}|^{2}\rangle$
@@
where the minus sign is due to fermions having exchange antisymmetry Griffiths and Schroeter (2018). By analogy of cold boson particles in a HBT interferometer, one-dimension spatial two-particle probability function is
@@
where $D$ is the length of S and single speed is assumed. The temporal two-particle probability function of fermions in a HBT interferometer is
@@
When two detectors are at the same space-time coordinates, the probability to detect two fermions equals 0. As the distance between these two detectors increases, the probability to detect two fermions increases to 1. There is two-particle antibunching effect for fermion particles in a HBT interferometer Henny et al. (1999); Jeltes et al. (2007), in contrast to the two-particle bunching effect for bosons in a HBT interferometer Jeltes et al. (2007); Brown and Twiss (1956b, a).
@@ VI.2.2 Second-order interference of two independent matter wave beams
Figure 18 shows the scheme for the second-order interference of two independent matter wave beams. S1 and S2 are two independent particle sources. BS is a particle beam splitter. D1 and D2 are two single-particle detectors. CC is two-particle coincidence count detection system.
First discuss the second-order interference of two single-particle sources. Assuming S1 and S2 are two identical single-particle sources, there are two different paths to trigger a two-particle coincidence count. The first one is the particle emitted by S1 is detected by D1 and the particle emitted by S2 is detected by D2. The probability is 1/2 and the two-particle probability amplitude is $A_{11,22}$. The second path is the particle emitted by S1 is detected by D2 and the particle emitted by S2 is detected by D1. The probability is 1/2 and the two-particle probability amplitude is $A_{12,21}$. If these different paths are indistinguishable, the two-particle probability function for bosons is
@@
$=$ $\langle|A_{11,22}+A_{12,21}|^{2}\rangle$
$=$ $\langle|e^{i\varphi_{1}}K_{11}e^{i\varphi_{2}}K_{22}+e^{i(\varphi_{1}+\frac{\pi}{2})}K_{12}e^{i(\varphi_{2}+\frac{\pi}{2})}K_{21}|^{2}\rangle$
@@
Substituting Eq. (104) into Eq. (116), one-dimension temporal two-particle probability function is
@@
where point particle source is assumed. It is impossible to detect two boson particles when these two detectors are at symmetrical positions. HOM dip with visibility higher than 50% can be observed Lopes et al. (2015).
With the same method, one-dimension temporal two-particle probability function for fermions is
@@
Then calculate the second-order interference of two independent and identical BECs in Fig. 18. There are three different ways to trigger a two-particle coincidence count. The first one is these two particles are both emitted by S1. The probability is 1/4 and there is one path, $A_{11,12}$. The second one is these two particles are both emitted by S2. The probability is 1/4 and there is one path, $A_{21,22}$. The third way is these two particles are emitted by S1 and S2 one each. The probability is 1/2 and there are two different paths, $A_{11,22}$ and $A_{12,21}$. When these four different paths are indistinguishable, the two-particle probability function is
@@
Substituting Eq. (104) into Eq. (VI.2.2), it is straightforward to have temporal two-particle probability function
@@
Compared to the second-order interference of two single boson sources, Eq. (117), the visibility of the second-order interference of two independent BECs can not exceed 50%, which is similar as the conclusion for the visibility of the second-order interference of two independent classical and nonclassical light beams Mandel (1983).
@@ VI.3 Third-order interference of matter wave
Based on the results above, it is easy to predict that there is three-particle bunching when a cold boson particle beams in a third-order HBT interferometer and no three-particle bunching for a BEC Burt et al. (1997). This section will concentrate on the third-order interference of fermions. The scheme for the third-order interference of one fermion beam is the same as the one shown in Fig. 27, in which the meanings of all the symbols are similar as the one for photons. There are six different ways to trigger a three-particle coincidence count and the corresponding three-particle probability amplitudes are $A_{a1,b2,c3}$, $A_{a1,b3,c2}$, $A_{a2,b1,c3}$, $A_{a2,b3,c1}$, $A_{a3,b1,c2}$ and $A_{a3,b2,c1}$. When these six different paths are indistinguishable, the three-particle probability function is
@@
$=$ $\langle|A_{a1,b2,c3}-A_{a1,b3,c2}-A_{a2,b1,c3}$
$+A_{a2,b3,c1}+A_{a3,b1,c2}-A_{a3,b2,c1}|^{2}\rangle,$
@@
Figure 35: Third-order temporal coherence function of fermions in a third-order HBT interferometer.
Figure 35 shows the normalized third-order temporal coherence function of fermions in a third-order HBT interferometer, which is proportional to Eq. (VI.3). The probability to detect three particles equals 0 when these three detectors are at the symmetrical positions, which is consistent with the results obtained by Liu et al. Liu (2016).
@@
VII Summary
In summary, we have presented a new quantum optical coherence theory based on Feynman’s path integral, which provides a useful tool for studying the physics of coherence besides classical optical coherence theory based on Maxwell’s electromagnetic theory and Glauber’s quantum optical coherence theory based on wave mechanics formalism of quantum mechanics. Some interesting conclusions are drawn. For instance, unlike the well-accepted conclusion that the electric field models are the same for thermal and laser light within the coherence time, it is predicted by quantum optical coherence theory based on Feynman’s path integral that there may be some difference between these two. The physics behind two-photon bunching of thermal light in a HBT interferometer and HOM dip of entangled photon pairs in a HOM interferometer is the same, which can be interpreted by two-photon interference. In order to show how quantum optical coherence theory based on Feynman’s path integral can help us to understand the physics of optical coherence better, subwavelength interference is taken as an example to be illustrated and it is shown that there is no intrinsic difference between the observed second-order subwavelength interference for two detectors scanning in the same direction and opposite directions. Finally, quantum particle coherence theory based on Feynman’s path integral is also given by analogy of the one in optics.
We believe quantum optical coherence theory based on Feynman’s path integral will be helpful to understand the physics of optical and atomic coherence in other fields, and eventually lead us to the answer to the question: what is photon?

### v2

@@ VI.1.2 First-order interference of two independent matter wave beams
Figure 34: Scheme for the first-order interference of two independent matter wave beams. S1 and S2 are two particle sources.
There are two different paths to trigger a particle detection event in the observation plane. The first one is that the detected particle is emitted S1. The probability is 1/2 and the probability amplitude is $A_{1x}$. The second one is that the detected particle is emitted S2. The probability is 1/2 and the probability amplitude is $A_{2x}$. If these two different paths are indistinguishable, the probability function of detecting a particle in the observation plane is
$P^{(1)}(x,t)$ $=$ $\langle|\frac{1}{\sqrt{2}}(A_{1x}+A_{2x})|^{2}\rangle$ (106)
$\propto$ $\langle|e^{i\varphi_{1j}}K_{1x}+e^{i\varphi_{2j}}K_{2x}|^{2}\rangle,$
where $\varphi_{1j}$ and $\varphi_{2j}$ are the initial phases of particles emitted by S1 and S2, respectively. With the same approximations above, the spatial probability function can be obtained via Eq. (106),
@@
The interference pattern is depend on the initial phases of particles. Comparing the experimental results of the degrees of second- and third-order coherence of cold atoms and BEC to the ones of thermal and laser light Yasuda and Shimizu (1996); Burt et al. (1997); Lopes et al. (2015), it is reasonable to assume that the coherence properties of cold atoms are similar as the ones of thermal light and the coherence properties of BEC are similar as the ones of laser light. The initial phases of particles in cold atoms are random. The initial phases of particles in one BEC are identical.
@@
where $N$ is the number of detected particles and $\varphi$ is a random phase due to the sum of $\varphi_{1j}-\varphi_{2j}$ for $N$ detected particles. The first-order interference pattern of two independent cold atom sources can not be observed.
@@
$P^{(1)}(x)$ $\propto$ $1+\cos[\frac{2\pi d}{L\lambda_{D}}x+(\varphi_{1j}-\varphi_{2j})].$ (109)
$\varphi_{1j}-\varphi_{2j}$ is a random but fixed number for two independent BECs. The first-order interference pattern can be observed Andrews et al. (1997), just like there is transient first-order interference pattern by superposing two independent laser light beams Magyar and Mandel (1963).
@@ VI.2.1 Second-order interference of one matter wave beam
The scheme for the second-order interference of one matter wave beam is similar as the one shown in Fig. 11 except that the source, S, is a particle source, BS is a beam splitter for particles and these two detectors are particle detectors.
Let us first discuss the second-order interference of one boson beam. Assuming S is a cold atom source, there are two different paths to trigger a two-particle coincidence count. The first path is that particle a is detected by D1 and particle b is detected by D2. The probability is 1/2 and the corresponding two-particle amplitude is $A_{a1,b2}$. The second path is that particle a is detected by D2 and particle b is detected by D1. The probability is 1/2 and the corresponding two-particle amplitude is $A_{a2,b1}$. If these two different paths are indistinguishable, the two-particle probability function is
@@
$=$ $\langle|\frac{1}{\sqrt{2}}(A_{a1,b2}+A_{a2,b1})|^{2}\rangle$
$\propto$ $\langle|e^{i\varphi_{a}}K_{a1}e^{i\varphi_{b}}K_{b2}+e^{i\varphi_{a}}K_{a2}e^{i\varphi_{b}}K_{b1}|^{2}\rangle$
@@
where $D$ is the length of S and single-speed particles are assumed.
The temporal two-particle probability function is
@@
When these two detectors are at the same space-time coordinate, the probability to detect two particles is two times of the one when these two detectors are in different coherence volumes. There is two-particle bunching for cold boson particles in a HBT interferometer Yasuda and Shimizu (1996), just like there is two-photon bunching for thermal light in a HBT interferometer.
Assuming S is a BEC in Fig. 11, there is only one path to trigger a two-particle coincidence count. The initial phases of the particles in a BEC are identical, just like the one for single-mode continuous-wave laser light. No two-particle bunching can be observed, which is consistent with Burt et al.’s experiments Burt et al. (1997).
Then discuss the second-order interference of one fermion beam. There are also two different paths to trigger a two-particle coincidence count and the corresponding two-particle probability amplitudes are $A_{a1,b2}$ and $A_{a2,b1}$. If these two different ways are indistinguishable, the two-particle probability function is
@@
$=$ $\langle|\frac{1}{\sqrt{2}}(A_{a1,b2}-A_{a2,b1})|^{2}\rangle$
$\propto$ $\langle|e^{i\varphi_{a}}K_{a1}e^{i\varphi_{b}}K_{b2}-e^{i\varphi_{a}}K_{a2}e^{i\varphi_{b}}K_{b1}|^{2}\rangle$
@@
where the minus sign is due to the exchange antisymmetry of fermions Griffiths and Schroeter (2018). By analogy of cold boson particles in a HBT interferometer, one-dimension spatial two-particle probability function for fermions in a HBT interferometer is
@@
where $D$ is the length of the source and single-speed particles are assumed.
The temporal two-particle probability function of fermions in a HBT interferometer is
@@
When these two detectors are at the same space-time coordinates, the probability of detecting two fermions equals 0. As the distance between these two detectors increases, the probability of detecting two fermions increases to 1. There is two-particle antibunching effect for fermions in a HBT interferometer Henny et al. (1999); Jeltes et al. (2007), which is in contrast to the two-particle bunching effect for bosons in a HBT interferometer Jeltes et al. (2007); Brown and Twiss (1956b, a).
@@ VI.2.2 Second-order interference of two independent matter wave beams
Figure 18 shows the scheme for the second-order interference of two independent matter wave beams. S1 and S2 are two independent particle sources. BS is a 1:1 particle beam splitter. D1 and D2 are two single-particle detectors. CC is two-particle coincidence count detection system.
First discuss the second-order interference of two single-boson sources. Assuming S1 and S2 are two identical and independent single-particle sources, there are two different paths to trigger a two-particle coincidence count. The first one is that the particle emitted by S1 is detected by D1 and the particle emitted by S2 is detected by D2. The probability is 1/2 and the two-particle probability amplitude is $A_{11,22}$. The second path is that the particle emitted by S1 is detected by D2 and the particle emitted by S2 is detected by D1. The probability is 1/2 and the two-particle probability amplitude is $A_{12,21}$. If these two different paths are indistinguishable, the two-particle probability function for bosons is
@@
$=$ $\langle\frac{1}{\sqrt{2}}(|A_{11,22}+A_{12,21})|^{2}\rangle$
$\propto$ $\langle|e^{i\varphi_{1}}K_{11}e^{i\varphi_{2}}K_{22}+e^{i(\varphi_{1}+\frac{\pi}{2})}K_{12}e^{i(\varphi_{2}+\frac{\pi}{2})}K_{21}|^{2}\rangle$
@@
where the extra $\pi/2$ phase is due to the reflection of BS comparing to the transmission like the one for photons. Substituting Eq. (104) into Eq. (116), temporal two-particle probability function is
@@
where point particle sources are assumed. It is impossible to detect two boson particles when these two detectors are at symmetrical positions of BS. Particle HOM dip with visibility higher than 50% was observed Lopes et al. (2015).
With the same method, temporal two-particle probability function in Fig. 18 for fermions is
@@
Then calculate the second-order interference of two independent and identical BECs in Fig. 18. There are three different ways to trigger a two-particle coincidence count. The first way is that these two particles are both emitted by S1. The probability is 1/4 and there is one path, $A_{11,12}$. The second one is that these two particles are both emitted by S2. The probability is 1/4 and there is one path, $A_{21,22}$. The third way is that these two particles are emitted by S1 and S2 one each. The probability is 1/2 and there are two different paths, $A_{11,22}$ and $A_{12,21}$. When these four different paths are indistinguishable, the two-particle probability function is
@@
Substituting Eq. (104) into Eq. (VI.2.2), it is straightforward to have temporal two-particle probability function,
@@
Comparing to the second-order interference of two single boson sources, the visibility of the second-order interference of two independent BECs can not exceed 50%, which is similar as the conclusion for the visibilities of the second-order interference of two independent classical and non-classical light beams Mandel (1983).
@@ VI.3 Third-order interference of matter wave
Based on the results above, it is easy to predict that there is three-particle bunching for a cold boson beam and no three-particle bunching for a BEC in a third-order HBT interferometer Burt et al. (1997). This section will focus on the third-order interference of fermions. The scheme for the third-order interference of one fermion beam is the same as the one shown in Fig. 27, in which the meanings of all the symbols are similar as the one for photons. There are six different ways to trigger a three-particle coincidence count and the corresponding three-particle probability amplitudes are $A_{a1,b2,c3}$, $A_{a1,b3,c2}$, $A_{a2,b1,c3}$, $A_{a2,b3,c1}$, $A_{a3,b1,c2}$ and $A_{a3,b2,c1}$. When these six different paths are indistinguishable, the three-particle probability function is
@@
$=$ $\langle|\frac{1}{\sqrt{6}}(A_{a1,b2,c3}-A_{a1,b3,c2}-A_{a2,b1,c3}$
$+A_{a2,b3,c1}+A_{a3,b1,c2}-A_{a3,b2,c1})|^{2}\rangle,$
@@
Figure 35: Third-order temporal coherence function of fermions in a third-order HBT interferometer. $g^{(3)}(t_{1},t_{2},t_{3})$ is the normalized third-order temporal coherence function. $t_{1}-t_{2}$ is the time difference between photon detection event at D1 and D2. Similar definition holds for $t_{2}-t_{3}$.
Figure 35 shows the normalized third-order temporal coherence function of fermions in a third-order HBT interferometer, which is proportional to Eq. (VI.3). The probability of detecting three particles when these three detectors are at the symmetrical positions equals 0, which is consistent with the results obtained by Liu et al. Liu (2016).
@@
VII Conclusions
It is interesting to notice that there is some difference between the points of view about the interference of light in classical optical coherence theory and quantum optical coherence theory based on Feynman’s path integral. In classical optical theory, the first-order interference of light is the foundation of the second- and higher-order interference of light. All the second- and higher-order interference of light is interpreted as the results of first-order interference of light. For instance, two-photon bunching of thermal light is interpreted as the intensity fluctuations correlation in classical optical coherence theory. In quantum optical coherence theory based on Feynman’s path integral, the first-, second-, and higher-order interference of light are interpreted by the same superposition principle. The first-order interference of light is not the foundation of second- and higher-order interference of light. There is no right and wrong about these two different points of view. It is just the interpretations about optical coherence are different in these two different theories.
In summary, we have presented a novel quantum optical coherence theory based on Feynman’s path integral, which provides a useful tool for studying the physics of optical coherence besides classical optical coherence theory based on Maxwell’s electromagnetic theory and Glauber’s quantum optical coherence theory based on matrix mechanics formalism of quantum mechanics. Some interesting conclusions are drawn. For instance, unlike the well-accepted conclusion that the electric field models are the same for thermal and laser light within the coherence time, it is predicted by quantum optical coherence theory based on Feynman’s path integral that there may be some difference between these two. The physics behind two-photon bunching of thermal light in a HBT interferometer and HOM dip of entangled photon pairs in a HOM interferometer is the same, both of which can be interpreted by two-photon interference. In order to show how quantum optical coherence theory based on Feynman’s path integral can be helpful to understand the physics of optical coherence, subwavelength interference is taken as an example and it is shown that there is no intrinsic difference between the observed two-photon subwavelength interference for two detectors scanning in the same direction and opposite directions. Finally, quantum particle coherence theory based on Feynman’s path integral is also given by analogy of the developed quantum optical coherence theory based on Feynman’s path integral.
We believe that quantum optical coherence theory based on Feynman’s path integral is helpful to understand the physics of optical and atomic coherence, and may eventually lead us to the answer to the question puzzling us for a very long time: what is light?